A club is creating a fund so that scholarships of $5000 can be paid out to the eligible students every year, for the next 20 years, starting next year. How much does the club need to invest today at 6.5% compounded annually?
Answer by Boreal(15235)


I had sinking fund by mistake
If use annuity formula,
P=payout(1-(1+r)^(-t))/r; r=0.065, t=20
=5000(1-(1.065^(-20))/0.065
=5000*0.7162/0.065
but round at end
=$55092.54